# Attention Entropy is a Key Factor: An Analysis of Parallel Context Encoding with Full-attention-based Pre-trained Language Models

### v1

@@ 1 Introduction
Figure 1: An overview. The upper part illustrates the full attention and parallel context encoding schemes (superscripts denote position encoding), while the lower part shows parallel encoding leads to irregularly high attention entropy for the query tokens. We explore two methods to reduce entropy (§4): “Sink” means adding attention sinks, and “SEL” means selective attention.
@@
2 Experimental Settings
@@ 2.1 Parallel Context Encoding
Table 1: Comparisons between full-attention and naive parallel encoding with Llama-3.1-8B (results are macro-averaged over all sub-tasks). Here, “P” indicates the parallel degree (the number of context sub-pieces). For each task, we also vary the total sequence lengths (considering 4K, 8K and 16K).
@@ 4.1.2 Selective Attention
  • Output. Finally, the mask $m$ is applied to the input probability tensor, and the final output attention probability tensor is obtained after a final re-normalization step to ensure that each row sums up to one.
Between the grouping and selection step, an optional reduction operation can be performed to aggregate information among tokens, heads or even layers. For example, if aggregating over the query-token dimension, $s_{group}$ is reduced from $[N_{layer},N_{head},L_{query},P]$ to $[N_{layer},N_{head},1,P]$. This reduction is useful in scenarios where the most relevant information comes from the same sub-piece for all tokens in the current query. If aggregating over heads, it can help to identify the most salient information-seeking head, such as the retrieval head (Wu et al., 2024). A more aggressive reduction can be performed across the first three dimensions, reducing the group score to $[1,1,1,P]$, which is exactly the same as a retrieval procedure. Notice that if we choose the layer dimension for aggregation, we need to forward the model twice since attention scores at later layers are not available when calculating previous layers; for other dimensions, the selective attention modification can be applied immediately after each attention score is calculated. We again use the sum of top-5 values as the reduction function to identify the most salient attention scores.
@@ 4.3.1 Variations on Selection Attention
We provide detailed ablation studies on various choices in the selection attention procedure. Since there are no clear query tokens in the LM task, our analysis primarily focuses on the other three tasks. We consider a typical scenario as the case study: using Llama-3.1-8B and 8K lengths, and adopting a difficult parallel degree of $P$=64. Firstly, we start with our default setting of aggregating over the Head and Token dimensions (denoted as “HT”), and vary the K value in sub-piece top-K selection process. The results indicate that the optimal setting varies by task: synthetic recall tasks benefit from a small K value, since they require precise information from a few pieces, while ICL and RAG perform better with slightly larger K values, since additional context information can be helpful. Next, we examine different ways of information aggregation (with TopK=5 for ICL and RAG, and TopK=2 for synthetic tasks). Once again, different tasks exhibit distinct patterns: aggregating over all layers, as in a retrieval setting, yields the best results for RAG, layer-wise selection is more effective for synthetic tasks, and query-level selection is not crucial for ICL. While a universal and consistent method that performs well across all tasks would be ideal, achieving this may be difficult and costly. One advantage of our method is its flexibility, allowing dynamic adjustment of configurations to suit the specific nature of each task.
@@ 5 Related Work
Since the introduction of self-attention in Transformers, analyzing the roles the attention mechanism plays has been a popular topic (Clark et al., 2019; Jain and Wallace, 2019; Serrano and Smith, 2019; Wiegreffe and Pinter, 2019; Bibal et al., 2022). The most relevant work to this study includes findings on attention sinks and specialized attention heads. Attention sinks refer to the initial tokens that attract most of the attention weights in many heads, and they have been utilized to extend LLMs to longer context lengths (Xiao et al., 2024; Han et al., 2024; Gu et al., 2024). These works also inspire our analyses on attention entropy and hidden state norms. Additionally, it has also been shown that there can be specialized attention heads that perform special functions, such as syntactic heads for encoding syntactical relations (Clark et al., 2019), retrieval heads for collecting information from long contexts (Wu et al., 2024), and induction heads that may constitute the mechanism for ICL Olsson et al. (2022). Our reduction operations in selective attention are also based on the hypothesis that there is a small portion of information-seeking heads that can collect the most salient features from the contexts.
@@ Appendix A Additional Results
  • Figure 8, 9, 10 and 11 illustrate more results of entropy reduction methods in different settings. The overall trends are similar to those in Figure 5.

### v2

@@ 1 Introduction
Figure 1: An overview. The upper sub-figures illustrate the full attention and parallel context encoding schemes, with superscripts indicating positional encoding. The lower sub-figure demonstrates that parallel encoding can result in irregularly high attention entropy for query tokens. In this example, the context is divided into four segments (P1–P4), which are individually encoded with the parallel encoding scheme (upper right). These pieces correspond to the sub-regions in the lower sub-figure. We explore two methods to reduce entropy (§4): “Sink” means adding attention sinks, and “SEL” means selective attention.
@@
2 Preliminaries and Settings
2.1 Attention Entropy
One central metric in our analysis is attention entropy, defined as $H(p)=-\sum_{i}p_{i}\cdot\log p_{i}$, where $p$ represents the attention probability distribution (attention weights). This metric quantifies the diversity of an attention head’s focus over previous contexts. A higher entropy value indicates that the query token distributes its attention broadly across many prior tokens, suggesting a more global contextual understanding. On the other hand, a lower entropy value reflects a more concentrated attention pattern, which is crucial for tasks requiring precise retrieval.
@@ 2.2 Parallel Context Encoding
Table 1: Comparisons between full-attention and naive parallel encoding with Llama-3.1-8B (results are macro-averaged over all sub-tasks). Here, “P” indicates the parallel degree (the number of context sub-pieces). For each task, we also vary the total sequence lengths (considering 4K, 8K and 16K). With larger “P” (larger than 10), the results are much worse than those of full attention.
@@ 4.1.2 Selective Attention
  • Output. Finally, the mask $m$ is applied to the input probability tensor (attention weights), and the final output attention probability tensor is obtained after a final re-normalization step to ensure that each row sums up to one.
Between the grouping and selection step, an optional reduction operation can be performed to aggregate information among tokens, heads or even layers. For example, if aggregating over the query-token dimension, $s_{group}$ is reduced from $[N_{layer},N_{head},L_{query},P]$ to $[N_{layer},N_{head},1,P]$. This reduction is useful in scenarios where the most relevant information comes from the same sub-piece for all tokens in the current query. If aggregating over heads, it can help to identify the most salient information-seeking head, such as the retrieval head (Wu et al., 2024). A more aggressive reduction can be performed across the first three dimensions, reducing the group score to $[1,1,1,P]$, which is exactly the same as a retrieval procedure. Notice that if we choose the layer dimension for aggregation, [^9] For other cases (without layer aggregation), this operation will bring only slight overhead, since we incrementally apply the selection mechanism for each layer (first performing softmax to obtain the attention scores, then performing selection, and finally re-normalizing). we need to forward the model twice since attention scores at later layers are not available when calculating previous layers; for other dimensions, the selective attention modification can be applied immediately after each attention score is calculated. We again use the sum of top-5 values as the reduction function to identify the most salient attention scores.
@@ 4.3.1 Variations on Selection Attention
We provide detailed ablation studies on various choices in the selection attention procedure. Since there are no clear query tokens in the LM task, our analysis primarily focuses on the other three tasks. We consider a typical scenario as the case study: using Llama-3.1-8B and 8K lengths, and adopting a difficult parallel degree of $P$=64. Firstly, we start with our default setting of aggregating over the Head and Token dimensions (denoted as “HT”), and vary the K value in sub-piece top-K selection process. The results indicate that the optimal setting varies by task: synthetic recall tasks benefit from a small K value, since they require precise information from a few pieces, while ICL [^10] Breaking down on ICL tasks, we further find some patterns: for example, for coarse-grained TREC (6 labels), the performance of “K=1” is closer to that of the best performing “K=5” (0.45 vs 0.48), while for other ICL tasks where there are tens of labels, “K=1” usually lags behind for more. and RAG perform better with slightly larger K values, since additional context information can be helpful. Next, we examine different ways of information aggregation (with TopK=5 for ICL and RAG, and TopK=2 for synthetic tasks). Once again, different tasks exhibit distinct patterns: aggregating over all layers, as in a retrieval setting, yields the best results for RAG, layer-wise selection is more effective for synthetic tasks, and query-level selection is not crucial for ICL. While a universal and consistent method that performs well across all tasks would be ideal, achieving this may be difficult and costly. One advantage of our method is its flexibility, allowing dynamic adjustment of configurations to suit the specific nature of each task.
@@ 5 Related Work
Since the introduction of self-attention in Transformers, analyzing the roles the attention mechanism plays has been a popular topic (Clark et al., 2019; Jain and Wallace, 2019; Serrano and Smith, 2019; Wiegreffe and Pinter, 2019; Bibal et al., 2022). The most relevant work to this study includes findings on attention sinks and specialized attention heads. Attention sinks refer to the initial tokens that attract most of the attention weights in many heads, and they have been utilized to extend LLMs to longer context lengths (Xiao et al., 2024; Han et al., 2024; Gu et al., 2024). These works also inspire our analyses on attention entropy and hidden state norms. Wang et al. (2024) also find that adding “sink” tokens improves performance from the perspective of precision. Additionally, it has also been shown that there can be specialized attention heads that perform special functions, such as syntactic heads for encoding syntactical relations (Clark et al., 2019), retrieval heads for collecting information from long contexts (Wu et al., 2024), and induction heads that may constitute the mechanism for ICL Olsson et al. (2022). Our reduction operations in selective attention are also based on the hypothesis that there is a small portion of information-seeking heads that can collect the most salient features from the contexts. Moreover, (Attanasio et al., 2022) propose Entropy-based Attention Regularization to mitigate bias by penalizing tokens with low self-attention entropy, demonstrating the usefulness of controlling attention entropy in a different scenario.
@@ Appendix A Additional Results
  • Figure 8, 9, 10, 11, 12 and 13 illustrate more results of entropy reduction methods in different settings. The overall trends are similar to those in Figure 5.
@@
Figure 10: The influence of the entropy reduction methods (with Llama-3.1-8B and 4K lengths). Notations are the same as those in Table 5.
Figure 11: The influence of the entropy reduction methods (with Llama-3.1-8B and 16K lengths). Notations are the same as those in Table 5.